### v1

# Risk-Neutral Pricing Model of Uniswap Liquidity Providing Position: A Stopping Time Approach

Abstract
In this paper, we introduce a novel pricing model for Uniswap V3, built upon stochastic processes and the Martingale Stopping Theorem. This model innovatively frames the valuation of positions within Uniswap V3. We further conduct a numerical analysis and examine the sensitivities through Greek risk measures to elucidate the model’s implications. The results underscore the model’s significant academic contribution and its practical applicability for Uniswap liquidity providers (LPs), particularly in assessing risk exposure and guiding hedging strategies.
Index Terms:
Uniswap, pricing model, greeks, perpetual option, risk management, hedging strategies
I Introduction
Uniswap[1], launched in November 2018, represents a decentralized, open-source protocol for cryptocurrency exchange. As the first automated market maker (AMM) to achieve widespread adoption, Uniswap enables users to conduct token swaps directly from their wallets via liquidity pools. In these pools, liquidity providers (LPs) earn fees by contributing their assets.
The introduction of Uniswap V3 brought significant advancements, most notably the concept of concentrated liquidity. This feature allows LPs to allocate their capital to specific price ranges, enhancing capital efficiency and yielding higher returns while simultaneously mitigating price risk. Furthermore, Uniswap v3 incorporates multiple fee tiers, thereby providing LPs with flexibility in selecting their compensation levels in accordance with their risk tolerance.
With the ongoing expansion of decentralized finance (DeFi) adoption, the demands on LPs in Uniswap v3 have intensified. LPs are now required to meticulously balance a myriad of factors, including risk, returns, and volatility, necessitating a sophisticated understanding and strategic approach to liquidity provision.
II Related Work and Limitations
II-A Loss-Versus-Rebalancing (LVR)
LVR is a metric designed to quantify the impact of adverse selection on automated market maker (AMM) LPs[2]. LVR captures the costs incurred by LPs due to stale prices that better-informed arbitrageurs exploit. Unlike impermanent loss, which only considers the starting and ending points of a trade, LVR compares the performance of LPs to a rebalancing portfolio, highlighting additional costs associated with providing liquidity.
LVR incorporates volatility into its calculations, making it a useful tool for selecting the optimal liquidity pool. However, it remains primarily a risk assessment tool. It does not evaluate the returns for LPs when exiting positions at different prices, nor does it provide guidance on the optimal range for market-making activities.
II-B Guillaume Lambert
Guillaume Lambert focused on pricing within the Black-Scholes-Merton (BSM) framework in his work[3][4]. He viewed liquidity positions as perpetual options and attempted to replicate liquidity positions with covered calls. He also extends his model towards implied volatility and hedging. However, a limitation of Lambert’s model is the specification of an expiration time T, which inherently makes it an approximation rather than a precise solution. Additionally, Lambert explored the use of stochastic processes to calculate the probability of prices remaining within a position range, which has greatly inspired our research. We aim to build upon this foundation by starting from stochastic processes to initiate our study.
II-C Martingale and first hitting problem with two boundaries
Unlike traditional derivatives with specific expiration dates, Uniswap V3 contracts are perpetual, persisting until the price breaches predefined boundaries or the contract holder voluntarily ceases market-making. In scenarios where the price crosses these preset boundaries, this constitutes a classic First Passage Problem, for which Darling and Siegert provided a derivation of the stopping time distribution using the Laplace Transform in 1953 [5]. We have further discovered that the expected value of the payoff under the risk-neutral measure contains exclusively terms of the form $\tau^{n}\exp(-r\tau)$ which are associated with the stopping time $\tau$. Therefore, we employ the Laplace Transform and its implications to derive the present value when the holder does not elect to terminate prematurely. In situations where the contract holder actively ceases market-making, we draw inspiration from the pricing methodology for Perpetual American Puts described by Shreve (2004) [6] and use numerical optimization to determine the optimal stopping strategy.
III Derivation
III-A Summary of Model Derivation
Uniswap V3 can be divided into two primary pricing components: the Liquidity Provider (LP) segment and the Rebate segment. The LP component is structured to accommodate either European or American option styles. Under the European construct, the Uniswap contract realizes its terminal value solely upon the price breaching the predetermined upper or lower thresholds. In contrast, the American framework allows the contract holder the flexibility to terminate the automated market-making process at any chosen point in time. The valuation of the Rebate segment hinges on the method selected by the contract holder for Rebate withdrawal. The upper limit scenario assumes the possibility of continuous Rebate extraction at the precise point of accrual. Meanwhile, the lower limit approach considers a singular, comprehensive withdrawal of all Rebates coinciding with the termination of the contract.
III-B Divide and Conquer
We initially assume that the ETHUSDT price follows a Geometric Brownian Motion (GBM) characterized by drift $\mu$ and annualized volatility $\sigma$ :
$dS(t)=\mu S(t)dt+\sigma S(t)dB_{t}$
or alternatively:
$S(t)=S_{0}\exp((\mu-\frac{\sigma^{2}}{2})t+\sigma B_{t})$
Define the unit price of ETHUSDT as follows
$P(t)=\frac{S(t)}{S_{0}}=\exp((\mu-\frac{\sigma^{2}}{2})t+\sigma B_{t})$
We partition the pricing structure of Uniswap V3 into two segments:
$V=V_{\text{LP}}+V_{\text{fee}}$
III-B1 Liquidity Provider segment
The unit price of the LP segment within the Uniswap V3 Contract, constrained by $S_{H}>S_{0}>S_{L}$, can be expressed as:
$V_{\text{LP}}(P_{t})$ =$=\begin{cases}L_{q}P(t)(\frac{1}{\sqrt{L}}-\frac{1}{\sqrt{H}}),P(t)<L\\ L_{q}(2\sqrt{P(t)}-\sqrt{L}-\frac{P(t)}{\sqrt{H}}),L<P(t)<H\\ L_{q}(\sqrt{H}-\sqrt{L}),H<P(t)\end{cases}$
where $H,L$ represent the unit bounds:
$H$ $=\frac{S_{H}}{S_{0}}$
$L$ $=\frac{S_{L}}{S_{0}}$
Additionally, the Liquidity Parameter is defined as:
$L_{q}=\frac{1}{2-\sqrt{L}-\frac{1}{\sqrt{H}}}$
III-B2 Rebate segment
The formulation for the rebate fee segment is comparatively straightforward:
$V_{\text{fee}}=C_{a}L_{q}t=365\times C_{d}L_{q}t$
where $C_{a},C_{d}$ are the daily and annual rebate rate.
III-C European LP Pricing
We are interested in evaluating:
$V(0)=\mathbb{E}(V(t))=\mathbb{E}(V_{\text{LP}}(t))+\mathbb{E}(V_{\text{fee}}(t))$
For the Uniswap V3 opsition, the stopping time $\tau$ is defined as:
$\tau=\inf\limits_{t<\infty}\{S(t)\leq S_{L}\text{ or }S(t)\geq S_{H}\}$
If the contract cannot be terminated before the stopping time, the current valuation of Uniswap V3 can be formulated as:
$V_{E}(0)$ $=\mathbb{E}(V(t))$
$=\mathbb{E}(V_{\text{LP}}(t))+\mathbb{E}(V_{\text{fee}}(t))$
$=\mathbb{E}[e^{-r\tau}\text{LP}_{H}|S(\tau)=S_{H}]+\mathbb{E}[e^{-r\tau}\text{LP}_{L}|S(\tau)=S_{L}]$
$\mathbb{+}E(V_{\text{fee}}(t))$
$=\text{LP}_{H}\mathbb{E}[e^{-r\tau}|S(\tau)=S_{H}]+\text{LP}_{L}\mathbb{E}[e^{-r\tau}|S(\tau)=S_{L}]$
$+\mathbb{E}(V_{\text{fee}}(t))$
The normalized log-unit price process of ETHUSDT is given by:
$\bar{W}(t)=\frac{\ln P(t)}{\sigma}=(\frac{\mu}{\sigma}-\frac{\sigma}{2})t+B_{t}$
Let us denote:
$\mu^{\prime}=\frac{\mu}{\sigma}-\frac{\sigma}{2}$
Thus, we have:
$\bar{W}(t)=x+\mu^{\prime}t+B_{t}$
where $x$ is the start point, $a,b$ are designated as the unit bounds:
$b$ $=\frac{\ln H}{\sigma},b^{\prime}=b-x$
$a$ $=\frac{\ln L}{\sigma},a^{\prime}=x-a$
Moreover, it also holds that:
$\tau=\inf\limits_{t<\infty}\{\bar{W}(t)\leq a\text{ or }\bar{W}(t)\geq b\}$
Utilizing the Laplace Transform, we can express the following expectations:
$\mathbb{E}[e^{-r\tau}|S(\tau)=S_{H}]=\exp({\mu^{\prime}b^{\prime}})\frac{\sinh[a^{\prime}\sqrt{\mu^{\prime 2}+2r}]}{\sinh[(b-a)\sqrt{\mu^{\prime 2}+2r}]}$
$\mathbb{E}[e^{-r\tau}|S(\tau)=S_{L}]=\exp({-\mu^{\prime}a^{\prime}})\frac{\sinh[(b^{\prime}\sqrt{\mu^{\prime 2}+2r}]}{\sinh[(b-a)\sqrt{\mu^{\prime 2}+2r}]}$
III-D American LP Pricing
In Euro-like pricing, liquidity providers can only exit when the price reaches the boundary. We further propose an American-like pricing model that allows liquidity providers to exit at any time. As Uniswap V3 contracts are perpetual, the termination of the contract relies exclusively on price movements, independent of time $t$. Therefore, it can be assumed that there exist boundaries determined solely by price levels. This implies that exit strategies should be based only on price considerations, analogous to the derivation of the perpetual American put option in Shreve’s work[6]. So we have:
$L<L_{1}<x<L_{2}<H$
where terminating automated market-making at prices $L_{1}$ or $L_{2}$. A set of stopping boundaries $L_{1},L_{2}$ can be determined through numerical optimization. First, we denote $c,d$ as previous section:
$d$ $=\frac{\ln L_{2}}{\sigma},d^{\prime}=d-x$
$c$ $=\frac{\ln L_{1}}{\sigma},c^{\prime}=x-c$
Then the equivalent price of the European contract:
$V_{E}$ $=LP_{L_{2}}\exp({\mu^{\prime}d^{\prime}})\frac{\sinh[c^{\prime}\sqrt{\mu^{\prime 2}+2r}]}{\sinh[(d-c)\sqrt{\mu^{\prime 2}+2r}]}$
$+LP_{L_{1}}\exp({-\mu^{\prime}c^{\prime}})\frac{\sinh[(d^{\prime}\sqrt{\mu^{\prime 2}+2r}]}{\sinh[(d-c)\sqrt{\mu^{\prime 2}+2r}]}$
The American contract can be priced using optimization techniques as following:
$V_{A}=\mathop{\max_{L<L_{1}<x<L_{2}<H}V_{E}(L_{1},L_{2})}$
III-E Rebate Pricing
For the rebate segment, if we assume that we can withdraw rebates continuously,(it is an upper bound of rebate)
[…]
$=C_{a}L_{q}\cdot\mathbb{E}\left[\frac{1-e^{-r\tau}}{r}\right]$
$=\frac{C_{a}L_{q}}{r}\cdot(1-\mathbb{E}\left[{e^{-r\tau}}\right])$
where,
$\mathbb{E}[e^{-r\tau}]:=F(r)$
$=\frac{e^{-\mu^{\prime}a^{\prime}}\operatorname{sh}\left(b^{\prime}\sqrt{2r+\mu^{\prime 2}}\right)+e^{\mu^{\prime}b^{\prime}}\operatorname{sh}\left(a^{\prime}\sqrt{2r+\mu^{\prime 2}}\right)}{\operatorname{sh}\left((b-a)\sqrt{2r+\mu^{\prime 2}}\right)}$
For the lower bound, we assume that we can withdraw all rebates when LPs close their position,
$\mathbb{E}(V_{\text{fee}}(t))=\mathbb{E}(C_{a}L_{q}e^{-r\tau}\tau)=C_{a}L_{q}\mathbb{E}(e^{-r\tau}\tau)$
where,
[…]
=$=\mathcal{L}\{\tau p(\tau)\}(r)$
$=-\frac{dF(r)}{dr}$
$=\frac{(b-a)(e^{-a^{\prime}\mu^{\prime}}\operatorname{sh}[b^{\prime}\sqrt{2r+\mu^{\prime 2}}]+e^{b^{\prime}\mu}\operatorname{sh}[a^{\prime}\sqrt{2r+\mu^{\prime 2}}])}{\operatorname{th}[(b-a)\sqrt{2r+\mu^{\prime 2}}]\operatorname{sh}[(b-a)\sqrt{2r+\mu^{\prime 2}}](\sqrt{2r+\mu^{\prime 2}})}$
$\ \ \ \ -\frac{b^{\prime}e^{-a^{\prime}\mu^{\prime}}\operatorname{ch}[b^{\prime}\sqrt{2r+\mu^{\prime 2}}]+ae^{b\mu}\operatorname{ch}[a^{\prime}\sqrt{2r+\mu^{\prime 2}}]}{\operatorname{sh}[(b-a)\sqrt{2r+\mu^{\prime 2}}](\sqrt{2r+\mu^{\prime 2}})}$
Figure 1: relation between spot price and present value
As illustrated in Figure 1, the blue line represents the variation of the payoff pricing model in response to changes in the spot price. The orange and green lines depict the corresponding numerical changes in the pricing model of a European option, where the orange line incorporates transaction costs and discounting, while the green line excludes these factors, both following the fluctuations in the spot price.We can observe that, under the current market parameters, if the option pricing model does not include transaction costs, it results in an expected loss at the current price, meaning that the present value is less than 1.Only by incorporating the potential transaction costs and their discounted value can we generate a positive return.
IV Sensitivity Analysis
In the previous sections, we derived the analytical solution of Uniswap V3 positions. In this section, we focus on the sensitivity analysis of the Uniswap V3 position, specifically examining the Greeks associated with a typical European option framework. We previously derived the closed-form solutions for the pricing mechanisms of Uniswap V3 positions. Here, we extend our analysis to assess how variations in underlying parameters affect the pricing—commonly referred to as ”Greeks.”
TABLE I: $C=0.2$ $r=0.05$ $\sigma=0.7$
Model PV Delta Gamma Vega Rho
Payoff 1.000 0.561 -0.632 nan nan
European 1.422 0.427 -0.747 1.720 -25.892
American 1.696 0.539 -0.989 0.843 -19.425
TABLE II: $C=0.04$ $r=0.05$ $\sigma=0.4$
Model PV Delta Gamma Vega Rho
Payoff 1.000 0.784 -0.883 nan nan
European 0.643 0.074 0.073 1.296 -7.546
American 1.018 0.765 -1.059 0.200 -2.019
Figure 2: Table I market condition pv plot
Figure 3: Table II market condition pv plot
IV-A European and American Present Value Comparison
We sampled two sets of market parameters based on actual market conditions to observe how the European and American pricing models reflect the actual value of Uniswap range order positions. As illustrated in Table II, Through numerical analysis, it can be observed that under the market parameters (C=0.04, r=0.05, $\sigma$=0.4), if we immediately establish a Uniswap range order position and exit the position only when it touches the boundary (i.e., behavior under European LP pricing), the present value of such a product will certainly be less than 1. In contrast, if we are able to exit at the optimal execution boundary before reaching the designated boundary, even without achieving a high value of C, we may still obtain a positive return. This is because the American LP pricing model indicates that if a better execution exit boundary exists outside the set LH, a positive return expectation is possible. Furthermore, when C increases (e.g., C=0.2, r=0.05, $\sigma$=0.7), As illustrated in Table I, even with a high value of $\sigma$ (which implies that we may reach the boundary and exit the position more quickly), the European LP pricing model still yields a positive expected value under the current spot price of 1. We also plotted the PV-spot graphs for both European and American pricing models under two sets of market parameters. Figures 2 and Figures 3 clearly demonstrate that, although the early exit boundary for the American option pricing model is narrower, it yields a significantly higher positive expected return compared to the European pricing model.
Figure 4: delta comparison between european and payoff
IV-B Delta Analysis
Delta refers to the sensitivity of the option’s price with respect to changes in the underlying asset’s price. As illustrated in Figure 4, the delta for European pricing and payoff valuation reveals noteworthy disparities due to inherent market parameters (C=0.05, r=0.04, $\sigma$=0.25). At the orange line’s zero point, corresponding to the range (1, 1.05) in the previous analysis, the delta suggests that during actual trading, the delta derived from payoff valuation might lead to residual risks characterized by the Greek value. This indicates potential risk exposure when delta hedging is performed under the assumption of risk neutrality.
Figure 5: relation between pv and $\sigma$ using different rebate calculation
IV-C The Impact of Volatility
The relationship between the volatility of the underlying asset and the pricing dynamics of the Uniswap V3 position is integral to our sensitivity analysis. Intuitively, we understand that according to the European option valuation model, the present value (PV) pricing should exhibit a monotonically negative correlation with volatility. This stems from the Automatic Market Maker (AMM) mechanism of Uniswap V3, which fundamentally behaves as a short volatility strategy. As the volatility increases, we anticipate a decrease in the valuation, signifying an inverse correlation.
However, the observed orange line in Figure 5deviates from this predicted relationship due to the dual pricing model for transaction fees, which has separate methodologies for calculating both upper and lower bounds. The rationale behind the lower bound is driven by the expectation of time until a transaction fee is collected, conceptualized through the theory of stopping times, which is directly proportional to volatility. Conversely, the upper bound assumes continuous trading throughout the transactional process, operating under the premise that serial transactions yield continuous fees, justifying the positive valuation associated with low volatility events.
We selected points with relatively low volatility, such as $\sigma$=0.02, to compare the pricing models of the two transaction fee calculation methods in terms of their present value (PV), as shown in Figure 6. In the PV-spot graph, the pricing model based on the upper bound of transaction fees consistently yields a higher value than the pricing model based on the lower bound of transaction fees.
Figure 6: relationship of pv-spot with different fee calculation
V future work
V-1 Hedging Strategies
Our pricing model is equipped to offer more accurate calculations of Greek risk measures such as delta, gamma, and vega. These calculations are crucial for developing sophisticated hedging strategies. By leveraging these measures, we can devise fully hedged strategies that maximize profitability while minimizing risk. Such strategies will not only benefit LPs by enhancing their returns and reducing risk but also contribute to the overall stability and efficiency of the DeFi ecosystem.
V-2 Adaptation to Uniswap v4
With the introduction of Uniswap v4, various new decentralized exchange (DEX) paradigms will be implemented, such as dynamic fee structures [7], which make $C$ a function of volatility. This simplification of the model allows LPs to adjust their positions less frequently in response to market conditions, making participation more accessible and efficient. Beyond dynamic fees, many potential enhancements and new models may emerge with Uniswap v4, requiring further adaptations and improvements to our frameworks.
V-3 Implied volatility
Our model is not only applicable for pricing but also serves as a tool for estimating implied volatility. Similar to the Black-Scholes-Merton (BSM) framework, we can infer the views of LPs on volatility through their positions. By aggregating these inferred volatilities, we can construct a liquidity surface that reflects the weighted perspectives of LPs.

### v2

@@
 The formulation for the rebate fee segment is comparatively straightforward:
-$V_{\text{fee}}=C_{a}L_{q}t=365\times C_{d}L_{q}t$
+$V_{\text{fee}}=C_{a}L_{q}V_{LP}(P_{t})t=365\times C_{d}L_{q}V_{LP}(P_{t})t$
 where $C_{a},C_{d}$ are the daily and annual rebate rate.
@@
 $=\mathbb{E}(V_{\text{LP}}(t))+\mathbb{E}(V_{\text{fee}}(t))$
-$=\mathbb{E}[e^{-r\tau}\text{LP}_{H}|S(\tau)=S_{H}]+\mathbb{E}[e^{-r\tau}\text{LP}_{L}|S(\tau)=S_{L}]$
-$\mathbb{+}E(V_{\text{fee}}(t))$
-$=\text{LP}_{H}\mathbb{E}[e^{-r\tau}|S(\tau)=S_{H}]+\text{LP}_{L}\mathbb{E}[e^{-r\tau}|S(\tau)=S_{L}]$
+$=\mathbb{E}[e^{-r\tau}V_{\text{LP}}(H)|S(\tau)=S_{H}]+\mathbb{E}[e^{-r\tau}V_{\text{LP}}(L)|S(\tau)=S_{L}]$
+$+\mathbb{E}(V_{\text{fee}}(t))$
+$=V_{\text{LP}}(H)\mathbb{E}[e^{-r\tau}|S(\tau)=S_{H}]+V_{\text{LP}}(L)\mathbb{E}[e^{-r\tau}|S(\tau)=S_{L}]$
 $+\mathbb{E}(V_{\text{fee}}(t))$
@@
 […]
-[…]
-$=C_{a}L_{q}\cdot\mathbb{E}\left[\frac{1-e^{-r\tau}}{r}\right]$
-$=\frac{C_{a}L_{q}}{r}\cdot(1-\mathbb{E}\left[{e^{-r\tau}}\right])$
+$=C_{a}L_{q}V_{LP}(P_{t})(\cdot\mathbb{E}\left[\int_{0}^{\tau}e^{-rt}dt\right]$
+$=C_{a}L_{q}V_{LP}(P_{t})\cdot\mathbb{E}\left[\frac{1-e^{-r\tau}}{r}\right]$
+$=\frac{C_{a}L_{q}V_{LP}(P_{t})}{r}\cdot(1-\mathbb{E}\left[{e^{-r\tau}}\right])$
 where,
@@
 For the lower bound, we assume that we can withdraw all rebates when LPs close their position,
-$\mathbb{E}(V_{\text{fee}}(t))=\mathbb{E}(C_{a}L_{q}e^{-r\tau}\tau)=C_{a}L_{q}\mathbb{E}(e^{-r\tau}\tau)$
+$\mathbb{E}(V_{\text{fee}}(t))=\mathbb{E}(C_{a}L_{q}V_{LP}(P_{t})e^{-r\tau}\tau)=C_{a}L_{q}V_{LP}(P_{t})\mathbb{E}(e^{-r\tau}\tau)$
 where,
@@
 Figure 1: Relation between spot price and present value
+Figure 2: Euro-Uniswap price distribution: Monte Carlo vs analytical solution
 As illustrated in Figure 1, the blue line represents the variation of the payoff pricing model in response to changes in the spot price.
-The orange and green lines depict the corresponding numerical changes in the pricing model of a European option, where the orange line incorporates transaction costs and discounting, while the green line excludes these factors, both following the fluctuations in the spot price.We can observe that, under the current market parameters, if the option pricing model does not include transaction costs, it results in an expected loss at the current price, meaning that the present value is less than 1.Only by incorporating the potential transaction costs and their discounted value can we generate a positive return.
+The orange and green lines depict the corresponding numerical changes in the pricing model of a European option, where the orange line incorporates transaction costs and discounting, while the green line excludes these factors, both following the fluctuations in the spot price.
+We can observe that, under the current market parameters, if the option pricing model does not include transaction costs, it results in an expected loss at the current price, meaning that the present value is less than 1.
+Only by incorporating the potential transaction costs and their discounted value can we generate a positive return.
+Figure 2 provides a Monte Carlo simulation to model the price dynamics of an asset following a Geometric Brownian Motion (GBM).
+The initial asset price $S_{0}$ was set to 1, with volatility $\sigma=0.6$, drift $\mu=0$, risk-free rate $r=0.04$, and boundary levels $H=1.2$ (upper bound) and $L=0.8$ (lower bound).
+We ran the simulation 10,000 times, calculating the payoff at each simulation based on whether the price reached the upper or lower boundary, then discounting the payoff to obtain the expected value.
+The Monte Carlo estimates closely matched the analytical solution derived from our new pricing formula, confirming the accuracy of our approach.
+The final distribution of the outcomes, represented by a histogram, exhibited a bimodal shape.
+This reflects the nature of the problem, where the asset price is either stopped at the upper bound $H$ or the lower bound $L$ due to the stopping conditions inherent in the Uniswap V3 liquidity model.
+The two peaks correspond to the outcomes where the price reaches these boundaries, capturing the distinctive characteristics of liquidity provision in a decentralized exchange framework.
+This bimodal distribution highlights the risk/reward trade-offs faced by liquidity providers in such a system, where extreme price movements lead to clear-cut outcomes based on boundary interactions.
+III-F Formulation of Case-Specific Models
+Based on the derivations from the foundational section, we have obtained the pricing formulas for the following four scenarios.
+For the pricing formula of the European LP option problem under a GBM process, assuming that transaction fees can only be collected upon exiting the position, the pricing formula is as follows:
+$V_{t}$
+$=LP_{H}e^{\mu^{\prime}b^{\prime}}\frac{\operatorname{sh}(a^{\prime}\sqrt{2r+\mu^{\prime 2}})}{\operatorname{sh}((b-a)\sqrt{2r+\mu^{\prime 2}})}$
+$+LP_{L}e^{\mu^{\prime}a^{\prime}}\frac{\operatorname{sh}(b^{\prime}\sqrt{2r+\mu^{\prime 2}})}{\operatorname{sh}((b-a)\sqrt{2r+\mu^{\prime 2}})}$
+$+C_{a}L_{q}V_{LP}(P_{t})$
+$(\frac{(b-a)(e^{\mu^{\prime}a^{\prime}}\operatorname{sh}(b^{\prime}\sqrt{2r+\mu^{\prime 2}})+e^{\mu^{\prime}b^{\prime}}\operatorname{sh}(a^{\prime}\sqrt{2r+\mu^{\prime 2}}))}{\sqrt{2r+\mu^{\prime 2}}\operatorname{th}((b-a)\sqrt{2r+\mu^{\prime 2}})\operatorname{sh}((b-a)\sqrt{2r+\mu^{\prime 2}})}$
+$-\frac{e^{\mu^{\prime}a^{\prime}}b^{\prime}\operatorname{ch}(b^{\prime}\sqrt{2r+\mu^{\prime 2}})+e^{\mu^{\prime}b^{\prime}}a^{\prime}\operatorname{ch}(a^{\prime}\sqrt{2r+\mu^{\prime 2}})}{\sqrt{2r+\mu^{\prime 2}}\operatorname{sh}((b-a)\sqrt{2r+\mu^{\prime 2}})})$
+When we assume that transaction fees can be continuously collected during the lifetime of the position, the pricing formula becomes:
+$V_{t}$ $=LP_{H}e^{\mu^{\prime}b^{\prime}}\frac{\operatorname{sh}(a^{\prime}\sqrt{2r+\mu^{\prime 2}})}{\operatorname{sh}((b-a)\sqrt{2r+\mu^{\prime 2}})}$
+$+LP_{L}e^{\mu^{\prime}a^{\prime}}\frac{\operatorname{sh}(b^{\prime}\sqrt{2r+\mu^{\prime 2}})}{\operatorname{sh}((b-a)\sqrt{2r+\mu^{\prime 2}})}$
+$+\frac{C_{a}L_{q}V_{LP}(P_{t})}{r}(1$
+$-\frac{e^{\mu^{\prime}a^{\prime}}\operatorname{sh}(b^{\prime}\sqrt{2r+\mu^{\prime 2}})+e^{\mu^{\prime}b^{\prime}}\operatorname{sh}(a^{\prime}\sqrt{2r+\mu^{\prime 2}})}{\operatorname{sh}((b-a)\sqrt{2r+\mu^{\prime 2}})})$
+For the pricing formula of the American LP option problem, assuming that transaction fees can only be collected upon exiting the position, the pricing formula is as follows:
+$V_{t}=LP_{L_{2}}e^{\mu^{\prime}d^{\prime}}\frac{\operatorname{sh}(c^{\prime}\sqrt{2r+\mu^{\prime 2}})}{\operatorname{sh}((d-c)\sqrt{2r+\mu^{\prime 2}})}$
+$+LP_{L_{1}}e^{\mu^{\prime}c^{\prime}}\frac{\operatorname{sh}(bd^{\prime}\sqrt{2r+\mu^{\prime 2}})}{\operatorname{sh}((d-c)\sqrt{2r+\mu^{\prime 2}})}+C_{a}L_{q}V_{LP}(P_{t})($
+$\frac{(d-c)(e^{\mu^{\prime}c^{\prime}}\operatorname{sh}(d^{\prime}\sqrt{2r+\mu^{\prime 2}})+e^{\mu^{\prime}d^{\prime}}\operatorname{sh}(c^{\prime}\sqrt{2r+\mu^{\prime 2}}))}{\sqrt{2r+\mu^{\prime 2}}\operatorname{th}((d-c)\sqrt{2r+\mu^{\prime 2}})\operatorname{sh}((d-c)\sqrt{2r+\mu^{\prime 2}})}$
+$-\frac{e^{\mu^{\prime}c^{\prime}}d^{\prime}\operatorname{ch}(d^{\prime}\sqrt{2r+\mu^{\prime 2}})+e^{\mu^{\prime}d^{\prime}}c^{\prime}\operatorname{ch}(c^{\prime}\sqrt{2r+\mu^{\prime 2}})}{\sqrt{2r+\mu^{\prime 2}}\operatorname{sh}((d-c)\sqrt{2r+\mu^{\prime 2}})})$
+When we assume that transaction fees can be continuously collected during the lifetime of the position, the pricing formula for the American LP option problem becomes:
+$V_{t}$ $=LP_{L_{2}}e^{\mu^{\prime}d^{\prime}}\frac{\operatorname{sh}(c^{\prime}\sqrt{2r+\mu^{\prime 2}})}{\operatorname{sh}((d-c)\sqrt{2r+\mu^{\prime 2}})}$
+$+LP_{L_{1}}e^{\mu^{\prime}c^{\prime}}\frac{\operatorname{sh}(bd^{\prime}\sqrt{2r+\mu^{\prime 2}})}{\operatorname{sh}((d-c)\sqrt{2r+\mu^{\prime 2}})}$
+$+\frac{C_{a}L_{q}V_{LP}(P_{t})}{r}(1$
+$-\frac{e^{\mu^{\prime}c^{\prime}}\operatorname{sh}(d^{\prime}\sqrt{2r+\mu^{\prime 2}})+e^{\mu^{\prime}d^{\prime}}\operatorname{sh}(c^{\prime}\sqrt{2r+\mu^{\prime 2}})}{\operatorname{sh}((d-c)\sqrt{2r+\mu^{\prime 2}})})$
 IV Sensitivity Analysis
@@
 American 1.018 0.765 -1.059 0.200 -2.019
-Figure 2: Table I market condition pv plot
-Figure 3: Table II market condition pv plot
-IV-A European and American Present Value Comparison
-We sampled two sets of market parameters based on actual market conditions to observe how the European and American pricing models reflect the actual value of Uniswap range order positions.
-As illustrated in Table II, Through numerical analysis, it can be observed that under the market parameters (C=0.04, r=0.05, $\sigma$=0.4), if we immediately establish a Uniswap range order position and exit the position only when it touches the boundary (i.e., behavior under European LP pricing), the present value of such a product will certainly be less than 1.
-In contrast, if we are able to exit at the optimal execution boundary before reaching the designated boundary, even without achieving a high value of C, we may still obtain a positive return.
-This is because the American LP pricing model indicates that if a better execution exit boundary exists outside the set LH, a positive return expectation is possible.
-Furthermore, when C increases (e.g., C=0.2, r=0.05, $\sigma$=0.7), As illustrated in Table I, even with a high value of $\sigma$ (which implies that we may reach the boundary and exit the position more quickly), the European LP pricing model still yields a positive expected value under the current spot price of 1.
-We also plotted the PV-spot graphs for both European and American pricing models under two sets of market parameters.
-Figures 2 and Figures 3 clearly demonstrate that, although the early exit boundary for the American option pricing model is narrower, it yields a significantly higher positive expected return compared to the European pricing model.
-Figure 4: delta comparison between european and payoff
-IV-B Delta Analysis
-Delta refers to the sensitivity of the option’s price with respect to changes in the underlying asset’s price.
-As illustrated in Figure 4, the delta for European pricing and payoff valuation reveals noteworthy disparities due to inherent market parameters (C=0.05, r=0.04, $\sigma$=0.25).
-At the orange line’s zero point, corresponding to the range (1, 1.05) in the previous analysis, the delta suggests that during actual trading, the delta derived from payoff valuation might lead to residual risks characterized by the Greek value.
-This indicates potential risk exposure when delta hedging is performed under the assumption of risk neutrality.
-Figure 5: relation between pv and $\sigma$ using different rebate calculation
-IV-C The Impact of Volatility
+Figure 3: European LP Pricing Analysis
+Our subsequent Greeks analysis will be conducted based on two sets of market parameters corresponding to different market conditions.
+One scenario features high transaction fees and high market volatility, referred to as ’good market situation’, as these conditions allow Uniswap V3 positions to accrue more fees.
+The other scenario involves low transaction fees and low market volatility, termed ’good market situation’.
+Taking the European LP option pricing model as an example, we present the valuation discrepancies between pricing under good/bad market conditions and the payoff model at different price levels.
+The good/bad market conditions referenced here are based on the market parameters used in Table I and Table I.
+The range order price interval is set at (0.8,1.2).
+As shown in Figure 3, under the calculations based on the European LP option pricing model, there is a significant difference between the current Uniswap V3 payoff value and its actual present value.
+This is because the current value of a Uniswap V3 position is not limited to the extractable payoff; it also incorporates the discounted value of future fees that can still be collected if the position is not immediately withdrawn.
+In scenarios with relatively low volatility and low transaction fees, the analysis shows that if the Uniswap V3 position is not withdrawn now, its future value will depreciate, and the discounted value of future fees will be insufficient to offset the loss caused by the position’s depreciation, as depicted by the green line in the Figure 3.
+Conversely, under good market conditions with higher volatility and transaction fees, the orange line in the Figure 3 demonstrates that the actual value of the Uniswap V3 position exceeds its payoff value.
+Figure 4: European LP Delta Analysis
+IV-A Delta Analysis
+In historical Uniswap V3 trading practices, to mitigate the uncertainty of future returns, delta hedging is often employed to maintain delta-neutral risk exposure for Uniswap V3 positions.
+However, based on our European LP option pricing model, we find that delta risk within the $(L,H)$ range differs from the delta derived from payoff-based calculations.
+Taking the comparison between the orange and blue lines in Figure 4 as an example (since it is uncommon to persist with Uniswap V3 positions under unfavorable market conditions), we observe that when the price is at 0.9, the delta value of the orange line exceeds that of the blue line.
+This indicates that the actual delta risk at this point is higher than the delta risk derived from payoff-based pricing.
+Consequently, more delta needs to be hedged to achieve delta neutrality.
+Conversely, when the price is at 1.1, the actual delta risk associated with holding the Uniswap V3 position is lower.
+This suggests that hedging excessive delta based on payoff-based calculations is unnecessary.
+Figure 5: European LP Gamma Analysis
+IV-B Gamma Analysis
+By conducting a gamma analysis of the European LP pricing model versus the payoff-based pricing model, we can observe from Figure 5 that holding a Uniswap V3 position inherently involves gamma risk.
+Gamma risk refers to the situation where large gamma values cause significant fluctuations in delta with small changes in the underlying price $S$.
+In such scenarios, the delta hedging strategy becomes less effective, as even minor changes in $S$ can lead to substantial delta risk, necessitating more frequent adjustments to maintain delta neutrality.
+Conversely, when gamma values are small, delta remains relatively stable, reducing the need for frequent hedging.
+As shown in the Figure 5, there is a significant difference between the gamma risk values derived from payoff-based modeling and those from the European LP pricing model.
+This highlights the higher requirements for managing gamma risk in Uniswap V3 positions, emphasizing the need for more refined risk control strategies.
+IV-C Vega Analysis
+Vega measures the sensitivity of the valuation of a Uniswap V3 European pricing position to changes in volatility.
+From the Figure 6, we observe the variation of vega with respect to spot prices under different market conditions.
+Notably, there is no vega estimation graph for the payoff-based pricing model, as this model assumes no correlation between price and volatility, and thus does not account for vega risk.
+From the Figure 6, it is evident that vega remains consistently negative within the defined $(L,H)$ range.
+This indicates that holding a Uniswap V3 position inherently represents a short volatility strategy, implying that lower volatility leads to higher returns.
+Figure 6: European LP Vega Analysis
+IV-D Greeks Comparison between different range order
+After analyzing the unique patterns exhibited by the Greeks under Uniswap V3 positions, we aim to compare the Greeks of Uniswap V3 positions established with different $(L,H)$ ranges to uncover the relationship between Uniswap V3 pricing and the choice of $(L,H)$.
+The results are illustrated in the figure.
+We set up three groups of Uniswap V3 range orders with price ranges of $(0.8,1.1)$, $(0.8,1.2)$, and $(0.8,1.3)$.
+The assumed market conditions are $r=0.05$, $C=0.2$, $\mu=0$, and $\sigma=0.7$.
+From Figure 7 ,we observe that as the pricing range widens, the European LP pricing model’s valuation results tend to increase.
+This is because a wider pricing range implies that the asset is expected to remain within the pricing range for a longer period, thereby extending the expected time for fee collection.
+Figure 7: Different Range LP Pricing Comparison
+An interesting observation arises from the delta comparison in Figure 8 .
+We find that as the pricing range widens, the delta values generally increase.
+This is because when the pricing range is very wide, the proportional impact of price fluctuations on the exchange between the underlying asset and currency diminishes relative to the total position value.
+Consequently, the position value remains strongly correlated with the underlying asset price, resulting in a higher delta.
+Conversely, when the pricing range is narrow, even small price movements can trigger significant exchange activity, weakening the correlation between position value and underlying asset price, thus reducing delta.
+Figure 8: Different Range LP Pricing delta Comparison
+Although wider pricing ranges lead to larger delta risks, the absolute value of gamma decreases, as shown in Figure 9.
+This indicates that delta hedging becomes easier for wider pricing ranges compared to narrower ones.
+With a broader pricing range, delta changes more smoothly as the spot price fluctuates, simplifying the delta hedging process.
+Figure 9: Different Range LP Pricing gamma Comparison
+Finally, as demonstrated in Figure 10, even with changes to the pricing range, vega remains strictly negative.
+This confirms that the relationship between price and volatility consistently exhibits a negative correlation.
+Figure 10: Different Range LP Pricing Pricing Analyiss versus vol
+To validate the effectiveness of our Greeks, we need to perform numerical tests. Specifically, based on the Taylor expansion method, the following equation holds:
+$\Delta V\approx\text{Delta}\times\Delta S+\frac{1}{2}\text{Gamma}\times\Delta S^{2}+\text{Vega}\times\Delta\sigma$
+We can use two sets of different parameters, $S$ and $\sigma$, to calculate the valuation change of our Uniswap V3 position before and after market fluctuations (based on the present value derived from our pricing model).
+If the Greeks we calculate, when integrated into the right-hand side of the equation, can estimate the pricing with an acceptable level of accuracy, it will sufficiently demonstrate the validity of our Greeks.
+The parameters we set are as follows, for the market before the fluctuation:
+$C=0.2,r=0.05,\mu=0,\sigma=0.7$
+$S=1,H=1.1,L=0.8$
+Taking ETH as an example, under general circumstances, we assume a 5 % decrease in price and a 10 % increase in volatility itself.
+Using our pricing model, the valuation of the Uniswap V3 position before and after the market fluctuation is calculated as follows:
+The Current Pv is: 1.02110809997664861.0211080999766486 1.02110809997664861.0211080999766486
+The new PV is: 0.99304363754849550.9930436375484955 0.99304363754849550.9930436375484955
+The Position Changing is $-0.028064462428153192$
+The Greeks obtained through calculation, along with the corresponding Taylor expansion results, are as follows:
+The Current delta is: 0.197906601375752440.19790660137575244 0.197906601375752440.19790660137575244
+The Current gamma is: $-11.101273578173476$
+The Current vega is: $-0.15958993211606654$
+Taylor Expansion estimation is: $-0.034943217289629125$
+The Current residential is: $-0.006878754861475933$
+Thus, under this level of market changes, the pricing estimation derived from the Taylor expansion exhibits an approximate error of 0.6% compared to the actual valuation of the Uniswap V3 position.
+IV-E The Impact of Volatility
 The relationship between the volatility of the underlying asset and the pricing dynamics of the Uniswap V3 position is integral to our sensitivity analysis.
@@
 In the PV-spot graph, the pricing model based on the upper bound of transaction fees consistently yields a higher value than the pricing model based on the lower bound of transaction fees.
+Figure 11: relation between pv and $\sigma$ using different rebate calculation
 Figure 12: relationship of pv-spot with different fee calculation

### v3

@@
 We further conduct a numerical analysis and examine the sensitivities through Greek risk measures to elucidate the model’s implications.
+Additionally, we demonstrate the model’s practical application by construct the implied volatility of the model and compare it to other exist model.
 The results underscore the model’s significant academic contribution and its practical applicability for Uniswap liquidity providers (LPs), particularly in assessing risk exposure and guiding hedging strategies.
 Index Terms:
-Uniswap, pricing model, greeks, perpetual option, risk management, hedging strategies
+Uniswap, pricing model, greeks, perpetual option, risk management, hedging strategies, implied volatility
 I Introduction
-Uniswap[1], launched in November 2018, represents a decentralized, open-source protocol for cryptocurrency exchange.
-As the first automated market maker (AMM) to achieve widespread adoption, Uniswap enables users to conduct token swaps directly from their wallets via liquidity pools.
-In these pools, liquidity providers (LPs) earn fees by contributing their assets.
-The introduction of Uniswap V3 brought significant advancements, most notably the concept of concentrated liquidity.
-This feature allows LPs to allocate their capital to specific price ranges, enhancing capital efficiency and yielding higher returns while simultaneously mitigating price risk.
-Furthermore, Uniswap V3 incorporates multiple fee tiers, thereby providing LPs with flexibility in selecting their compensation levels in accordance with their risk tolerance.
-With the ongoing expansion of decentralized finance (DeFi) adoption, the demands on LPs in Uniswap v3 have intensified.
-LPs are now required to meticulously balance a myriad of factors, including risk, returns, and volatility, necessitating a sophisticated understanding and strategic approach to liquidity provision.
-II Related Work and Limitations
-II-A Loss-Versus-Rebalancing (LVR)
-LVR is a metric designed to quantify the impact of adverse selection on automated market maker (AMM) LPs[2].
-LVR captures the costs incurred by LPs due to stale prices that better-informed arbitrageurs exploit.
-Unlike impermanent loss, which only considers the starting and ending points of a trade, LVR compares the performance of LPs to a rebalancing portfolio, highlighting additional costs associated with providing liquidity.
-LVR incorporates volatility into its calculations, making it a useful tool for selecting the optimal liquidity pool.
-However, it remains primarily a risk assessment tool.
-It does not evaluate the returns for LPs when exiting positions at different prices, nor does it provide guidance on the optimal range for market-making activities.
-II-B Guillaume Lambert
-Guillaume Lambert focused on pricing within the Black-Scholes-Merton (BSM) framework in his work[3][4].
-He viewed liquidity positions as perpetual options and attempted to replicate liquidity positions with covered calls.
-He also extends his model towards implied volatility and hedging.
-However, a limitation of Lambert’s model is the specification of an expiration time T, which inherently makes it an approximation rather than a precise solution.
-Additionally, Lambert explored the use of stochastic processes to calculate the probability of prices remaining within a position range, which has greatly inspired our research.
-We aim to build upon this foundation by starting from stochastic processes to initiate our study.
-II-C Martingale and first hitting problem with two boundaries
-Unlike traditional derivatives with specific expiration dates, Uniswap V3 contracts are perpetual, persisting until the price breaches predefined boundaries or the contract holder voluntarily ceases market-making.
-In scenarios where the price crosses these preset boundaries, this constitutes a classic First Passage Problem, for which Darling and Siegert provided a derivation of the stopping time distribution using the Laplace Transform in 1953 [5].
-We have further discovered that the expected value of the payoff under the risk-neutral measure contains exclusively terms of the form $\tau^{n}\exp(-r\tau)$ which are associated with the stopping time $\tau$.
-Therefore, we employ the Laplace Transform and its implications to derive the present value when the holder does not elect to terminate prematurely.
-In situations where the contract holder actively ceases market-making, we draw inspiration from the pricing methodology for Perpetual American Puts described by Shreve (2004) [6] and use numerical optimization to determine the optimal stopping strategy.
-III Derivation
-III-A Summary of Model Derivation
-Uniswap V3 can be divided into two primary pricing components: the Liquidity Provider (LP) segment and the Rebate segment.
-The LP component is structured to accommodate either European or American option styles.
+I-A Background and Motivation
+Uniswap, launched in 2018, is one of the leading decentralized exchanges (DEXs) in the cryptocurrency market, enabling users to swap tokens directly from their wallets through liquidity pools, without relying on traditional intermediaries.
+As the first automated market maker (AMM)[1] to achieve widespread adoption, Uniswap has played a pivotal role in the growth of decentralized finance (DeFi).
+The introduction of Uniswap V3 marked a significant advancement in decentralized exchange technology, most notably through the concept of concentrated liquidity.
+This feature allows LPs to allocate capital to specific price ranges, improving capital efficiency and enhancing potential returns while simultaneously reducing price risk.
+With the ongoing growth of DeFi adoption, LPs now face a more complex landscape, balancing a range of factors such as risk, returns, and volatility.
+This requires a sophisticated understanding of liquidity provision strategies.
+In this work, we treat Uniswap V3 as an exotic perpetual derivative and use the Black-Scholes framework, along with the Martingale Stopping Theorem, to derive a closed-form pricing formula.
+Additionally, we perform a Monte Carlo simulation and sensitivity analysis to evaluate the model’s performance.
+We also compare our model with other state-of-the-art models to assess its advantages and limitations.
+Looking ahead, Uniswap V4 is expected to introduce even more advanced features that will further enhance the efficiency and flexibility of decentralized exchange mechanisms.
+Our model is designed to incorporate these new features of Uniswap V4, ensuring its adaptability to future developments.
+I-B Related Work
+Our work builds on and extends two broad strands of literature: (i) DeFi research on AMMs, and (ii) traditional financial engineering studies on market-making.
+Below, we discuss the most relevant works and highlight how our research contributes to these fields.
+I-B1 Automated Market Makers
+A significant advancement in AMM research is the concept of Loss-Versus-Rebalancing (LVR), introduced by [2] to quantify the adverse selection costs faced by LPs in AMM.
+LVR measures the performance gap between an LP’s portfolio and a dynamically rebalanced portfolio, capturing the losses incurred due to stale prices exploited by arbitrageurs.
+Another important contribution is the work of [3], who modeled Uniswap V3 liquidity positions as perpetual options within the Black-Scholes-Merton (BSM) framework.
+Lambert’s work treats LP positions as covered calls and explores implied volatility and hedging strategies.
+Recent studies have further advanced the understanding of AMMs.
+For example, [4] analyze the tradeoff between risks and returns for LPs in Uniswap V3, while [5] study the predictable losses of LPs in a continuous-time.
+[6] explore the hedging of impermanent losses.
+Closest to our work are the models in [7] and [8], which focus on fee revenue and use approximation techniques to obtain dynamic strategies.
+I-B2 Market-Making in Traditional Finance
+Our work is also related to the market-making literature in traditional finance.
+Early works include [9] and [10], which model market-making as an inventory management problem and derive optimal bid-ask spread strategies.
+These foundational studies have been extended in many directions; see [11] and [12] for recent developments
+Another particularly relevant area is the study of perpetual derivatives, which share similarities with the perpetual nature of Uniswap V3 liquidity positions.
+[13] provides a comprehensive treatment of perpetual put options using stochastic processes, deriving the distribution of stopping times and optimal stopping strategies.
+I-C Contributions
+Our work is structured into four key steps, each addressing a critical aspect of pricing Uniswap V3 liquidity positions and extending the model’s applicability:
+Step 1: Perpetual Liquidity Position Modeling
+We model Uniswap V3 liquidity positions as perpetual derivatives, using the Martingale Stopping Theorem and Laplace Transform to derive closed-form pricing formulas.
+This approach captures the perpetual nature of liquidity positions and eliminates the need for fixed expiration times.
+Step 2: Monte Carlo Validation
+We validate our closed-form formulas through Monte Carlo simulations. The results align closely with the analytical solutions, confirming the accuracy and robustness of our model.
+Step 3: Greeks and Implied Volatility
+We conduct a comprehensive analysis of Greek risk measures (Delta, Gamma, Vega) and derive implied volatility from our model.
+A key comparison with the LVR framework highlights the advantages of our approach in capturing implied volatility.
+Step 4: Generalization to Other AMMs and Uniswap V4
+We discuss the potential extension of our model to other AMMs and its adaptation to the dynamic fee structures expected in Uniswap V4.
+While our current focus is on Uniswap V3, we outline possible modifications to the framework to accommodate future developments in AMM design.
+II Derivation
+II-A Summary of Model Derivation
+Uniswap V3 can be divided into two primary pricing components: the LPs segment and the Rebate segment.
+The LPs component is structured to accommodate either European or American option styles.
 Under the European construct, the Uniswap contract realizes its terminal value solely upon the price breaching the predetermined upper or lower thresholds.
@@
 Meanwhile, the lower limit approach considers a singular, comprehensive withdrawal of all Rebates coinciding with the termination of the contract.
-III-B Divide and Conquer
+II-B Divide and Conquer
 We initially assume that the ETHUSDT price follows a Geometric Brownian Motion (GBM) characterized by drift $\mu$ and annualized volatility $\sigma$ :
@@
 $V=V_{\text{LP}}+V_{\text{fee}}$
-III-B1 Liquidity Provider segment
-The unit price of the LP segment within the Uniswap V3 Contract, constrained by $S_{H}>S_{0}>S_{L}$, can be expressed as:
+II-B1 Liquidity Provider segment
+The unit price of the LPs segment within the Uniswap V3 Contract, constrained by $S_{H}>S_{0}>S_{L}$, can be expressed as:
 $V_{\text{LP}}(P_{t})$ =$=\begin{cases}L_{q}P(t)(\frac{1}{\sqrt{L}}-\frac{1}{\sqrt{H}}),P(t)<L\\ L_{q}(2\sqrt{P(t)}-\sqrt{L}-\frac{P(t)}{\sqrt{H}}),L<P(t)<H\\ L_{q}(\sqrt{H}-\sqrt{L}),H<P(t)\end{cases}$
@@
 $L_{q}=\frac{1}{2-\sqrt{L}-\frac{1}{\sqrt{H}}}$
-III-B2 Rebate segment
+II-B2 Rebate segment
 The formulation for the rebate fee segment is comparatively straightforward:
@@
 where $C_{a},C_{d}$ are the daily and annual rebate rate.
-III-C European LP Pricing
+II-C European LP Pricing
 We are interested in evaluating:
@@
 $\mathbb{E}[e^{-r\tau}|S(\tau)=S_{L}]=\exp({-\mu^{\prime}a^{\prime}})\frac{\sinh[(b^{\prime}\sqrt{\mu^{\prime 2}+2r}]}{\sinh[(b-a)\sqrt{\mu^{\prime 2}+2r}]}$
-III-D American LP Pricing
+II-D American LP Pricing
 In Euro-like pricing, liquidity providers can only exit when the price reaches the boundary.
@@
 Therefore, it can be assumed that there exist boundaries determined solely by price levels.
-This implies that exit strategies should be based only on price considerations, analogous to the derivation of the perpetual American put option in Shreve’s work[6].
+This implies that exit strategies should be based only on price considerations, analogous to the derivation of the perpetual American put option in [13]’s work.
 So we have:
@@
 $V_{A}=\mathop{\max_{L<L_{1}<x<L_{2}<H}V_{E}(L_{1},L_{2})}$
-III-E Rebate Pricing
+II-E Rebate Pricing
 For the rebate segment, if we assume that we can withdraw rebates continuously,(it is an upper bound of rebate)
 […]
-$=C_{a}L_{q}V_{LP}(P_{t})(\cdot\mathbb{E}\left[\int_{0}^{\tau}e^{-rt}dt\right]$
-$=C_{a}L_{q}V_{LP}(P_{t})\cdot\mathbb{E}\left[\frac{1-e^{-r\tau}}{r}\right]$
-$=\frac{C_{a}L_{q}V_{LP}(P_{t})}{r}\cdot(1-\mathbb{E}\left[{e^{-r\tau}}\right])$
+[…]
+$=C_{a}L_{q}\cdot\mathbb{E}\left[\frac{1-e^{-r\tau}}{r}\right]$
+$=\frac{C_{a}L_{q}}{r}\cdot(1-\mathbb{E}\left[{e^{-r\tau}}\right])$
 where,
@@
 For the lower bound, we assume that we can withdraw all rebates when LPs close their position,
-$\mathbb{E}(V_{\text{fee}}(t))=\mathbb{E}(C_{a}L_{q}V_{LP}(P_{t})e^{-r\tau}\tau)=C_{a}L_{q}V_{LP}(P_{t})\mathbb{E}(e^{-r\tau}\tau)$
+$\mathbb{E}(V_{\text{fee}}(t))=\mathbb{E}(C_{a}L_{q}e^{-r\tau}\tau)=C_{a}L_{q}\mathbb{E}(e^{-r\tau}\tau)$
 where,
@@
 As illustrated in Figure 1, the blue line represents the variation of the payoff pricing model in response to changes in the spot price.
-The orange and green lines depict the corresponding numerical changes in the pricing model of a European option, where the orange line incorporates transaction costs and discounting, while the green line excludes these factors, both following the fluctuations in the spot price.
-We can observe that, under the current market parameters, if the option pricing model does not include transaction costs, it results in an expected loss at the current price, meaning that the present value is less than 1.
-Only by incorporating the potential transaction costs and their discounted value can we generate a positive return.
-Figure 2 provides a Monte Carlo simulation to model the price dynamics of an asset following a Geometric Brownian Motion (GBM).
-The initial asset price $S_{0}$ was set to 1, with volatility $\sigma=0.6$, drift $\mu=0$, risk-free rate $r=0.04$, and boundary levels $H=1.2$ (upper bound) and $L=0.8$ (lower bound).
-We ran the simulation 10,000 times, calculating the payoff at each simulation based on whether the price reached the upper or lower boundary, then discounting the payoff to obtain the expected value.
-The Monte Carlo estimates closely matched the analytical solution derived from our new pricing formula, confirming the accuracy of our approach.
-The final distribution of the outcomes, represented by a histogram, exhibited a bimodal shape.
-This reflects the nature of the problem, where the asset price is either stopped at the upper bound $H$ or the lower bound $L$ due to the stopping conditions inherent in the Uniswap V3 liquidity model.
-The two peaks correspond to the outcomes where the price reaches these boundaries, capturing the distinctive characteristics of liquidity provision in a decentralized exchange framework.
-This bimodal distribution highlights the risk/reward trade-offs faced by liquidity providers in such a system, where extreme price movements lead to clear-cut outcomes based on boundary interactions.
-III-F Formulation of Case-Specific Models
+The orange and green lines depict the corresponding numerical changes in the pricing model of a European option, where the orange line incorporates transaction costs and discounting, while the green line excludes these factors, both following the fluctuations in the spot price.We can observe that, under the current market parameters, if the option pricing model does not include transaction costs, it results in an expected loss at the current price, meaning that the present value is less than 1.Only by incorporating the potential transaction costs and their discounted value can we generate a positive return.
+Figure 2 shows a Monte Carlo simulation of asset price dynamics under GBM, with $S_{0}=1,\sigma=0.6,\mu=0,r=0.04,H=1.2,L=0.8$.Running 10,000 simulations, we calculated payoffs based on boundary hits, discounted them, and matched the results to our analytical solution, confirming accuracy.
+The outcome histogram exhibits a bimodal shape, reflecting price stops at $H$ or $L$ due to Uniswap V3’s stopping conditions.
+The peaks correspond to boundary interactions, capturing the risk/reward trade-offs for liquidity providers in decentralized exchanges.
+II-F Formulation of Case-Specific Models
 Based on the derivations from the foundational section, we have obtained the pricing formulas for the following four scenarios.
-For the pricing formula of the European LP option problem under a GBM process, assuming that transaction fees can only be collected upon exiting the position, the pricing formula is as follows:
+For the pricing formula of the European LPs option problem under a GBM process, assuming that transaction fees can only be collected upon exiting the position, the pricing formula is as follows:
 $V_{t}$
-$=LP_{H}e^{\mu^{\prime}b^{\prime}}\frac{\operatorname{sh}(a^{\prime}\sqrt{2r+\mu^{\prime 2}})}{\operatorname{sh}((b-a)\sqrt{2r+\mu^{\prime 2}})}$
-$+LP_{L}e^{\mu^{\prime}a^{\prime}}\frac{\operatorname{sh}(b^{\prime}\sqrt{2r+\mu^{\prime 2}})}{\operatorname{sh}((b-a)\sqrt{2r+\mu^{\prime 2}})}$
-$+C_{a}L_{q}V_{LP}(P_{t})$
-$(\frac{(b-a)(e^{\mu^{\prime}a^{\prime}}\operatorname{sh}(b^{\prime}\sqrt{2r+\mu^{\prime 2}})+e^{\mu^{\prime}b^{\prime}}\operatorname{sh}(a^{\prime}\sqrt{2r+\mu^{\prime 2}}))}{\sqrt{2r+\mu^{\prime 2}}\operatorname{th}((b-a)\sqrt{2r+\mu^{\prime 2}})\operatorname{sh}((b-a)\sqrt{2r+\mu^{\prime 2}})}$
-$-\frac{e^{\mu^{\prime}a^{\prime}}b^{\prime}\operatorname{ch}(b^{\prime}\sqrt{2r+\mu^{\prime 2}})+e^{\mu^{\prime}b^{\prime}}a^{\prime}\operatorname{ch}(a^{\prime}\sqrt{2r+\mu^{\prime 2}})}{\sqrt{2r+\mu^{\prime 2}}\operatorname{sh}((b-a)\sqrt{2r+\mu^{\prime 2}})})$
+$=LP_{H}e^{\mu^{\prime}b^{\prime}}\frac{sh(a^{\prime}\sqrt{2r+\mu^{\prime 2}})}{sh((b-a)\sqrt{2r+\mu^{\prime 2}})}$
+$+LP_{L}e^{\mu^{\prime}a^{\prime}}\frac{sh(b^{\prime}\sqrt{2r+\mu^{\prime 2}})}{sh((b-a)\sqrt{2r+\mu^{\prime 2}})}$
+$+C_{a}L_{q}*$
+$(\frac{(b-a)(e^{\mu^{\prime}a^{\prime}}sh(b^{\prime}\sqrt{2r+\mu^{\prime 2}})+e^{\mu^{\prime}b^{\prime}}sh(a^{\prime}\sqrt{2r+\mu^{\prime 2}}))}{\sqrt{2r+\mu^{\prime 2}}th((b-a)\sqrt{2r+\mu^{\prime 2}})sh((b-a)\sqrt{2r+\mu^{\prime 2}})}$
+$-\frac{e^{\mu^{\prime}a^{\prime}}b^{\prime}ch(b^{\prime}\sqrt{2r+\mu^{\prime 2}})+e^{\mu^{\prime}b^{\prime}}a^{\prime}ch(a^{\prime}\sqrt{2r+\mu^{\prime 2}})}{\sqrt{2r+\mu^{\prime 2}}sh((b-a)\sqrt{2r+\mu^{\prime 2}})})$
 When we assume that transaction fees can be continuously collected during the lifetime of the position, the pricing formula becomes:
-$V_{t}$ $=LP_{H}e^{\mu^{\prime}b^{\prime}}\frac{\operatorname{sh}(a^{\prime}\sqrt{2r+\mu^{\prime 2}})}{\operatorname{sh}((b-a)\sqrt{2r+\mu^{\prime 2}})}$
-$+LP_{L}e^{\mu^{\prime}a^{\prime}}\frac{\operatorname{sh}(b^{\prime}\sqrt{2r+\mu^{\prime 2}})}{\operatorname{sh}((b-a)\sqrt{2r+\mu^{\prime 2}})}$
-$+\frac{C_{a}L_{q}V_{LP}(P_{t})}{r}(1$
-$-\frac{e^{\mu^{\prime}a^{\prime}}\operatorname{sh}(b^{\prime}\sqrt{2r+\mu^{\prime 2}})+e^{\mu^{\prime}b^{\prime}}\operatorname{sh}(a^{\prime}\sqrt{2r+\mu^{\prime 2}})}{\operatorname{sh}((b-a)\sqrt{2r+\mu^{\prime 2}})})$
-For the pricing formula of the American LP option problem, assuming that transaction fees can only be collected upon exiting the position, the pricing formula is as follows:
-$V_{t}=LP_{L_{2}}e^{\mu^{\prime}d^{\prime}}\frac{\operatorname{sh}(c^{\prime}\sqrt{2r+\mu^{\prime 2}})}{\operatorname{sh}((d-c)\sqrt{2r+\mu^{\prime 2}})}$
-$+LP_{L_{1}}e^{\mu^{\prime}c^{\prime}}\frac{\operatorname{sh}(bd^{\prime}\sqrt{2r+\mu^{\prime 2}})}{\operatorname{sh}((d-c)\sqrt{2r+\mu^{\prime 2}})}+C_{a}L_{q}V_{LP}(P_{t})($
-$\frac{(d-c)(e^{\mu^{\prime}c^{\prime}}\operatorname{sh}(d^{\prime}\sqrt{2r+\mu^{\prime 2}})+e^{\mu^{\prime}d^{\prime}}\operatorname{sh}(c^{\prime}\sqrt{2r+\mu^{\prime 2}}))}{\sqrt{2r+\mu^{\prime 2}}\operatorname{th}((d-c)\sqrt{2r+\mu^{\prime 2}})\operatorname{sh}((d-c)\sqrt{2r+\mu^{\prime 2}})}$
-$-\frac{e^{\mu^{\prime}c^{\prime}}d^{\prime}\operatorname{ch}(d^{\prime}\sqrt{2r+\mu^{\prime 2}})+e^{\mu^{\prime}d^{\prime}}c^{\prime}\operatorname{ch}(c^{\prime}\sqrt{2r+\mu^{\prime 2}})}{\sqrt{2r+\mu^{\prime 2}}\operatorname{sh}((d-c)\sqrt{2r+\mu^{\prime 2}})})$
-When we assume that transaction fees can be continuously collected during the lifetime of the position, the pricing formula for the American LP option problem becomes:
-$V_{t}$ $=LP_{L_{2}}e^{\mu^{\prime}d^{\prime}}\frac{\operatorname{sh}(c^{\prime}\sqrt{2r+\mu^{\prime 2}})}{\operatorname{sh}((d-c)\sqrt{2r+\mu^{\prime 2}})}$
-$+LP_{L_{1}}e^{\mu^{\prime}c^{\prime}}\frac{\operatorname{sh}(bd^{\prime}\sqrt{2r+\mu^{\prime 2}})}{\operatorname{sh}((d-c)\sqrt{2r+\mu^{\prime 2}})}$
-$+\frac{C_{a}L_{q}V_{LP}(P_{t})}{r}(1$
-$-\frac{e^{\mu^{\prime}c^{\prime}}\operatorname{sh}(d^{\prime}\sqrt{2r+\mu^{\prime 2}})+e^{\mu^{\prime}d^{\prime}}\operatorname{sh}(c^{\prime}\sqrt{2r+\mu^{\prime 2}})}{\operatorname{sh}((d-c)\sqrt{2r+\mu^{\prime 2}})})$
-IV Sensitivity Analysis
+$V_{t}$ $=LP_{H}e^{\mu^{\prime}b^{\prime}}\frac{sh(a^{\prime}\sqrt{2r+\mu^{\prime 2}})}{sh((b-a)\sqrt{2r+\mu^{\prime 2}})}$
+$+LP_{L}e^{\mu^{\prime}a^{\prime}}\frac{sh(b^{\prime}\sqrt{2r+\mu^{\prime 2}})}{sh((b-a)\sqrt{2r+\mu^{\prime 2}})}$
+$+\frac{C_{a}L_{q}}{r}(1$
+$-\frac{e^{\mu^{\prime}a^{\prime}}sh(b^{\prime}\sqrt{2r+\mu^{\prime 2}})+e^{\mu^{\prime}b^{\prime}}sh(a^{\prime}\sqrt{2r+\mu^{\prime 2}})}{sh((b-a)\sqrt{2r+\mu^{\prime 2}})})$
+For the pricing formula of the American LPs option problem, assuming that transaction fees can only be collected upon exiting the position, the pricing formula is as follows:
+$V_{t}=LP_{L_{2}}e^{\mu^{\prime}d^{\prime}}\frac{sh(c^{\prime}\sqrt{2r+\mu^{\prime 2}})}{sh((d-c)\sqrt{2r+\mu^{\prime 2}})}$
+$+LP_{L_{1}}e^{\mu^{\prime}c^{\prime}}\frac{sh(bd^{\prime}\sqrt{2r+\mu^{\prime 2}})}{sh((d-c)\sqrt{2r+\mu^{\prime 2}})}$
+$+C_{a}L_{q}(\frac{(d-c)(e^{\mu^{\prime}c^{\prime}}sh(d^{\prime}\sqrt{2r+\mu^{\prime 2}})+e^{\mu^{\prime}d^{\prime}}sh(c^{\prime}\sqrt{2r+\mu^{\prime 2}}))}{\sqrt{2r+\mu^{\prime 2}}th((d-c)\sqrt{2r+\mu^{\prime 2}})sh((d-c)\sqrt{2r+\mu^{\prime 2}})}$
+$-\frac{e^{\mu^{\prime}c^{\prime}}d^{\prime}ch(d^{\prime}\sqrt{2r+\mu^{\prime 2}})+e^{\mu^{\prime}d^{\prime}}c^{\prime}ch(c^{\prime}\sqrt{2r+\mu^{\prime 2}})}{\sqrt{2r+\mu^{\prime 2}}sh((d-c)\sqrt{2r+\mu^{\prime 2}})})$
+When we assume that transaction fees can be continuously collected during the lifetime of the position, the pricing formula for the American LPs option problem becomes:
+$V_{t}$ $=LP_{L_{2}}e^{\mu^{\prime}d^{\prime}}\frac{sh(c^{\prime}\sqrt{2r+\mu^{\prime 2}})}{sh((d-c)\sqrt{2r+\mu^{\prime 2}})}$
+$+LP_{L_{1}}e^{\mu^{\prime}c^{\prime}}\frac{sh(bd^{\prime}\sqrt{2r+\mu^{\prime 2}})}{sh((d-c)\sqrt{2r+\mu^{\prime 2}})}$
+$+\frac{C_{a}L_{q}}{r}(1$
+$-\frac{e^{\mu^{\prime}c^{\prime}}sh(d^{\prime}\sqrt{2r+\mu^{\prime 2}})+e^{\mu^{\prime}d^{\prime}}sh(c^{\prime}\sqrt{2r+\mu^{\prime 2}})}{sh((d-c)\sqrt{2r+\mu^{\prime 2}})})$
+III Sensitivity Analysis
 In the previous sections, we derived the analytical solution of Uniswap V3 positions.
@@
 Figure 3: European LP Pricing Analysis
-Our subsequent Greeks analysis will be conducted based on two sets of market parameters corresponding to different market conditions.
-One scenario features high transaction fees and high market volatility, referred to as ’good market situation’, as these conditions allow Uniswap V3 positions to accrue more fees.
-The other scenario involves low transaction fees and low market volatility, termed ’good market situation’.
-Taking the European LP option pricing model as an example, we present the valuation discrepancies between pricing under good/bad market conditions and the payoff model at different price levels.
-The good/bad market conditions referenced here are based on the market parameters used in Table I and Table I.
-The range order price interval is set at (0.8,1.2).
-As shown in Figure 3, under the calculations based on the European LP option pricing model, there is a significant difference between the current Uniswap V3 payoff value and its actual present value.
-This is because the current value of a Uniswap V3 position is not limited to the extractable payoff; it also incorporates the discounted value of future fees that can still be collected if the position is not immediately withdrawn.
-In scenarios with relatively low volatility and low transaction fees, the analysis shows that if the Uniswap V3 position is not withdrawn now, its future value will depreciate, and the discounted value of future fees will be insufficient to offset the loss caused by the position’s depreciation, as depicted by the green line in the Figure 3.
-Conversely, under good market conditions with higher volatility and transaction fees, the orange line in the Figure 3 demonstrates that the actual value of the Uniswap V3 position exceeds its payoff value.
+Our Greeks analysis will use two market parameter sets: one with high fees and high volatility (”good market”), and another with low fees and low volatility (”bad market”).
+Using the European LP option pricing model, we compare valuation differences between good/bad market conditions and payoff models across price levels, with a range order interval of (0.8, 1.2).
+As shown in Figure 3, the European LP model reveals a significant gap between Uniswap V3’s current payoff and its present value, as the latter includes discounted future fees.
+In low volatility and fee scenarios, not withdrawing the position leads to value depreciation, with future fees failing to offset losses (green line).
+Conversely, in high volatility and fee ”good market” conditions, the position’s actual value exceeds its payoff (orange line).
 Figure 4: European LP Delta Analysis
-IV-A Delta Analysis
-In historical Uniswap V3 trading practices, to mitigate the uncertainty of future returns, delta hedging is often employed to maintain delta-neutral risk exposure for Uniswap V3 positions.
-However, based on our European LP option pricing model, we find that delta risk within the $(L,H)$ range differs from the delta derived from payoff-based calculations.
-Taking the comparison between the orange and blue lines in Figure 4 as an example (since it is uncommon to persist with Uniswap V3 positions under unfavorable market conditions), we observe that when the price is at 0.9, the delta value of the orange line exceeds that of the blue line.
-This indicates that the actual delta risk at this point is higher than the delta risk derived from payoff-based pricing.
-Consequently, more delta needs to be hedged to achieve delta neutrality.
-Conversely, when the price is at 1.1, the actual delta risk associated with holding the Uniswap V3 position is lower.
-This suggests that hedging excessive delta based on payoff-based calculations is unnecessary.
+III-A Delta Analysis
+In Uniswap V3 trading, delta hedging is often used to manage delta-neutral risk.
+However, our European LP option pricing model reveals that delta risk within $(L,H)$ the range differs from payoff-based delta.
+For example, in Figure 4, the orange line (model delta) exceeds the blue line (payoff-based delta) at a price of 0.9, indicating higher actual delta risk and requiring more hedging.
+Conversely, at 1.1, model delta risk is lower, making excessive hedging unnecessary.
 Figure 5: European LP Gamma Analysis
-IV-B Gamma Analysis
-By conducting a gamma analysis of the European LP pricing model versus the payoff-based pricing model, we can observe from Figure 5 that holding a Uniswap V3 position inherently involves gamma risk.
-Gamma risk refers to the situation where large gamma values cause significant fluctuations in delta with small changes in the underlying price $S$.
-In such scenarios, the delta hedging strategy becomes less effective, as even minor changes in $S$ can lead to substantial delta risk, necessitating more frequent adjustments to maintain delta neutrality.
-Conversely, when gamma values are small, delta remains relatively stable, reducing the need for frequent hedging.
-As shown in the Figure 5, there is a significant difference between the gamma risk values derived from payoff-based modeling and those from the European LP pricing model.
-This highlights the higher requirements for managing gamma risk in Uniswap V3 positions, emphasizing the need for more refined risk control strategies.
-IV-C Vega Analysis
+III-B Gamma Analysis
+Gamma analysis of the European LP pricing model versus the payoff-based model (Figure 5) shows that Uniswap V3 positions inherently carry gamma risk.
+Large gamma values cause delta to fluctuate significantly with small changes in $S$, reducing delta hedging effectiveness and requiring frequent adjustments.
+Conversely, small gamma values stabilize delta, minimizing hedging needs.
+Figure 5 reveals a significant gap between gamma risks from payoff-based and European LP models, highlighting the need for more refined risk control in Uniswap V3 positions.
+III-C Vega Analysis
 Vega measures the sensitivity of the valuation of a Uniswap V3 European pricing position to changes in volatility.
@@
 Figure 6: European LP Vega Analysis
-IV-D Greeks Comparison between different range order
-After analyzing the unique patterns exhibited by the Greeks under Uniswap V3 positions, we aim to compare the Greeks of Uniswap V3 positions established with different $(L,H)$ ranges to uncover the relationship between Uniswap V3 pricing and the choice of $(L,H)$.
-The results are illustrated in the figure.
-We set up three groups of Uniswap V3 range orders with price ranges of $(0.8,1.1)$, $(0.8,1.2)$, and $(0.8,1.3)$.
-The assumed market conditions are $r=0.05$, $C=0.2$, $\mu=0$, and $\sigma=0.7$.
-From Figure 7 ,we observe that as the pricing range widens, the European LP pricing model’s valuation results tend to increase.
-This is because a wider pricing range implies that the asset is expected to remain within the pricing range for a longer period, thereby extending the expected time for fee collection.
+III-D Greeks Comparison between different range order
+[…]
 Figure 7: Different Range LP Pricing Comparison
-An interesting observation arises from the delta comparison in Figure 8 .
-We find that as the pricing range widens, the delta values generally increase.
-This is because when the pricing range is very wide, the proportional impact of price fluctuations on the exchange between the underlying asset and currency diminishes relative to the total position value.
-Consequently, the position value remains strongly correlated with the underlying asset price, resulting in a higher delta.
-Conversely, when the pricing range is narrow, even small price movements can trigger significant exchange activity, weakening the correlation between position value and underlying asset price, thus reducing delta.
 Figure 8: Different Range LP Pricing delta Comparison
-Although wider pricing ranges lead to larger delta risks, the absolute value of gamma decreases, as shown in Figure 9.
-This indicates that delta hedging becomes easier for wider pricing ranges compared to narrower ones.
-With a broader pricing range, delta changes more smoothly as the spot price fluctuates, simplifying the delta hedging process.
 Figure 9: Different Range LP Pricing gamma Comparison
-Finally, as demonstrated in Figure 10, even with changes to the pricing range, vega remains strictly negative.
-This confirms that the relationship between price and volatility consistently exhibits a negative correlation.
 Figure 10: Different Range LP Pricing Pricing Analyiss versus vol
-To validate the effectiveness of our Greeks, we need to perform numerical tests. Specifically, based on the Taylor expansion method, the following equation holds:
-$\Delta V\approx\text{Delta}\times\Delta S+\frac{1}{2}\text{Gamma}\times\Delta S^{2}+\text{Vega}\times\Delta\sigma$
-We can use two sets of different parameters, $S$ and $\sigma$, to calculate the valuation change of our Uniswap V3 position before and after market fluctuations (based on the present value derived from our pricing model).
-If the Greeks we calculate, when integrated into the right-hand side of the equation, can estimate the pricing with an acceptable level of accuracy, it will sufficiently demonstrate the validity of our Greeks.
-The parameters we set are as follows, for the market before the fluctuation:
-$C=0.2,r=0.05,\mu=0,\sigma=0.7$
-$S=1,H=1.1,L=0.8$
-Taking ETH as an example, under general circumstances, we assume a 5 % decrease in price and a 10 % increase in volatility itself.
-Using our pricing model, the valuation of the Uniswap V3 position before and after the market fluctuation is calculated as follows:
-The Current Pv is: 1.02110809997664861.0211080999766486 1.02110809997664861.0211080999766486
-The new PV is: 0.99304363754849550.9930436375484955 0.99304363754849550.9930436375484955
-The Position Changing is $-0.028064462428153192$
-The Greeks obtained through calculation, along with the corresponding Taylor expansion results, are as follows:
-The Current delta is: 0.197906601375752440.19790660137575244 0.197906601375752440.19790660137575244
-The Current gamma is: $-11.101273578173476$
-The Current vega is: $-0.15958993211606654$
-Taylor Expansion estimation is: $-0.034943217289629125$
-The Current residential is: $-0.006878754861475933$
-Thus, under this level of market changes, the pricing estimation derived from the Taylor expansion exhibits an approximate error of 0.6% compared to the actual valuation of the Uniswap V3 position.
-IV-E The Impact of Volatility
-The relationship between the volatility of the underlying asset and the pricing dynamics of the Uniswap V3 position is integral to our sensitivity analysis.
-Intuitively, we understand that according to the European option valuation model, the present value (PV) pricing should exhibit a monotonically negative correlation with volatility.
-This stems from the Automatic Market Maker (AMM) mechanism of Uniswap V3, which fundamentally behaves as a short volatility strategy.
-As the volatility increases, we anticipate a decrease in the valuation, signifying an inverse correlation.
-However, the observed orange line in Figure 11deviates from this predicted relationship due to the dual pricing model for transaction fees, which has separate methodologies for calculating both upper and lower bounds.
-The rationale behind the lower bound is driven by the expectation of time until a transaction fee is collected, conceptualized through the theory of stopping times, which is directly proportional to volatility.
-Conversely, the upper bound assumes continuous trading throughout the transactional process, operating under the premise that serial transactions yield continuous fees, justifying the positive valuation associated with low volatility events.
-We selected points with relatively low volatility, such as $\sigma$=0.02, to compare the pricing models of the two transaction fee calculation methods in terms of their present value (PV), as shown in Figure 12.
-In the PV-spot graph, the pricing model based on the upper bound of transaction fees consistently yields a higher value than the pricing model based on the lower bound of transaction fees.
+From Figure 7, wider ranges increase European LP valuations, as assets stay longer within the range, extending fee collection time.
+Figure 8 shows delta rises with range width, as price fluctuations impact exchanges less relative to total position value, strengthening price correlation.
+Conversely, narrow ranges weaken this correlation, reducing delta.
+Figure 9 indicates that while wider ranges increase delta risk, gamma decreases, making delta hedging easier due to smoother delta changes.
+Finally, Figure 10 confirms vega remains negative across ranges, highlighting the negative price-volatility relationship.
+III-E The Impact of Volatility
+The relationship between underlying asset volatility and Uniswap V3 pricing is key to our sensitivity analysis.
+Intuitively, the European option model suggests PV should negatively correlate with volatility, as Uniswap V3’s AMM acts as a short volatility strategy.
+Higher volatility thus lowers valuation.
+However, Figure 11’s orange line deviates due to the dual pricing model for fees.
+The lower bound reflects fee collection time, proportional to volatility via stopping time theory.
+The upper bound assumes continuous trading, justifying positive valuation in low volatility scenarios.
+Figure 12 compares PV for $\sigma=0.02$. The upper-bound fee model consistently yields higher PV than the lower-bound model.
 Figure 11: relation between pv and $\sigma$ using different rebate calculation
 Figure 12: relationship of pv-spot with different fee calculation
-V future work
-V-1 Hedging Strategies
-Our pricing model is equipped to offer more accurate calculations of Greek risk measures such as delta, gamma, and vega.
-These calculations are crucial for developing sophisticated hedging strategies.
-By leveraging these measures, we can devise fully hedged strategies that maximize profitability while minimizing risk.
-Such strategies will not only benefit LPs by enhancing their returns and reducing risk but also contribute to the overall stability and efficiency of the DeFi ecosystem.
-V-2 Adaptation to Uniswap v4
-With the introduction of Uniswap v4, various new decentralized exchange (DEX) paradigms will be implemented, such as dynamic fee structures [7], which make $C$ a function of volatility.
-This simplification of the model allows LPs to adjust their positions less frequently in response to market conditions, making participation more accessible and efficient.
-Beyond dynamic fees, many potential enhancements and new models may emerge with Uniswap v4, requiring further adaptations and improvements to our frameworks.
-V-3 Implied volatility
-Our model is not only applicable for pricing but also serves as a tool for estimating implied volatility.
-Similar to the Black-Scholes-Merton (BSM) framework, we can infer the views of LPs on volatility through their positions.
-By aggregating these inferred volatilities, we can construct a liquidity surface that reflects the weighted perspectives of LPs.
+IV Application of model
+In this section, we apply model to calculate the implied volatility and try to generalize the model to more specific AMM case.
+IV-A Implied Volatility
+Building upon the foundation of pricing research, we can also deduce the implied volatility of corresponding Uni V3 positions by inversely solving the pricing model.We can aggregate the implied volatility views of all Uni V3 investors at a cross-section, weighting them to observe the historical performance of the weighted implied volatility views.
+Notably, when we compare the historical performance of implied volatility with LVR model, we uncover some fascinating patterns.
+The volatility inferred by LVR is based on the assumption of no arbitrage.
+We possess an instantaneous LVR formula, and we also understand that the collection of fees is approximately a linear model over an extremely short period.
+According to the [14], under the no-arbitrage assumption, the instantaneous LVR should equate to the instantaneous fees collected.
+$C\lambda=\frac{\lambda\sigma^{2}}{4}$
+Based on this established equality, the sigma term within the instantaneous LVR formula can be inversely calculated by other parameters in the equation, thereby yielding an implied volatility formula based on LVR.
+$\sigma=2\sqrt{C}$
+We collected historical transaction data for the ETH-USDC-005 trading pair on the Ethereum blockchain from June to December 2024 and created a comparative chart of implied volatility, the figure 13.
+We can observe that the historical performance of the weighted IV largely aligns with the trend of the IV inferred by LVR.
+Additionally, the graph presents unweighted and non-aggregated IV samples, revealing that the IV deduced by LVR predominantly resides at the upper echelon of the IV samples.
+This phenomenon can be attributed to our model’s assumption that future returns are subject to discounting.
+Consequently, our model inherently presumes the advantage of early boundary attainment for profit realization.
+As a result, the IV derived from our model is comparatively underestimated relative to the IV inferred by LVR.
+Figure 13: IV histogram Comparison
+While the implied volatility expressions are notably similar across the three models, key differences lie in their assumptions and specifics:
+  1. 1.
+    Lambert’s model assumes theta approximates a Dirac delta as time approaches zero, deriving the final option price.
+LVR also relies on infinitesimal time.
+Our model uses stopping time of a stochastic process to determine expiration, providing expected maturity time.
+  2. 2.
+    LVR and Lambert assume zero risk-free rate, equating future earnings to present value. Our model incorporates the risk-free rate for more accurate present value estimation of Uni V3 positions.
+  3. 3.
+    LVR and Lambert can only give the instantaneous value of Uni V3 position, in hence they can’t capture greeks. On the contrary, our model can easily calculate greeks to measure the risk.
+  4. 4.
+    LVR and Lambert focus on liquidity pool pricing, ignoring range order differences.
+We argue range orders reflect investor views on price and volatility, making price range relevant to Uni V3 pricing
+  5. 5.
+    LVR and Lambert assume zero-drift geometric Brownian motion. Our risk-neutral model allows flexible drift assumptions.
+  6. 6.
+    Like LVR, our model excels in scalability, as demonstrated in the next subsection.
+We summarize three models as following table, RNP represent our model’s name risk-neutral-pricing model.
+RNP Lambert LVR
+Maturity stopping time 0 flexible
+risk free flexible 0 0
+greeks support non-support non-support
+AMM range relevant irrelevant irrelevant
+price process GBM 0 drift GBM 0 drift GBM
+scalability other AMM none other AMM
+IV-B Extend to other AMM case
+IV-B1 Example of application in Uniswap V2
+In the case of Uniswap V2, as the liquidity bounds (LH) approach zero and infinity, the stopping time for the euro-situation tends toward infinity, rendering the model ineffective.
+However, the modeling approach for the amer-situation can still be applied to Uniswap V2.
+This is because, as the duration of the position extends, the discounted value of the collected fees has a diminishing impact on overall returns, making fee collection less attractive to investors.
+The price range at which investors choose to exit their Uniswap V2 positions influences the expected duration of the position.
+Therefore, based on the price range at which investors actively close their positions, our model can still price the Uniswap V2 position.
+As the formula represent:
+$L_{1}<x<L_{2}$
+$d=\frac{\ln L_{2}}{\sigma},d^{\prime}=d-x$
+$c=\frac{\ln L_{1}}{\sigma},c^{\prime}=x-c$
+So we have the formula:
+$V_{E}$ $=LP_{L_{2}}^{V2\ case}\exp({\mu^{\prime}d^{\prime}})\frac{\sinh[c^{\prime}\sqrt{\mu^{\prime 2}+2r}]}{\sinh[(d-c)\sqrt{\mu^{\prime 2}+2r}]}$
+$+LP_{L_{1}}^{V2\ case}\exp({-\mu^{\prime}c^{\prime}})\frac{\sinh[(d^{\prime}\sqrt{\mu^{\prime 2}+2r}]}{\sinh[(d-c)\sqrt{\mu^{\prime 2}+2r}]}$
+This formula can solve the Uniswap V2 pricing problem.
+IV-B2 Application in Uniswap V4: Dynamic fee example
+Uniswap V4 has a unique mechanism called dynamic fee.
+Unlike V3’s constant fee.
+Dynamic fee can be setting more flexible.
+One kind of dynamic fee can be adjusted in real-time based on various market conditions, like volatility.
+It means we can regard fee as a variable that is directly influenced by volatility.
+Our model can easily adapt dynamic fee mechanism, by adjusting the internal parameter C of the model into a function of volatility.
+We believe our model will also be highly adaptable to the pricing of Uniswap V4 positions for variety hooks.
+V Conclusion
+In this paper, we have constructed a risk-neutral pricing model based on the stopping time theorem to price Uniswap V3 positions.
+Our model capture the value of each investor’s range order is predicated on the discount of all potential earnings during the period until they first exit the position.
+From this hypothesis, we regard the Uniswap V3 position as an option pricing problem and have derived several analytical solution based on euro-situation and amer-situation, along with the corresponding Greeks based on these pricing formulas.
+The Greeks can be utilized to measure the risk exposure faced by investors holding this position.
+We have also employed MC simulations to verify the accuracy of our formula.
+This pricing model successfully delineates the relationship between the value of a Uniswap V3 position, its range setting, and the market’s volatility.
+Through this model, investors can discern the implied volatility perspective underlying their Uni V3 positions and compare it with the prevailing liquidity views in the market.
+Moreover, with an accurate estimation of risk exposure, investors can more readily manage their risks and steer clear of undesired exposures.